What are equilibrium price and quantity in the Chinese coal market if the demand and supply of Chinese coal are
Qd = 60 − 0.5P.
Qs = −2 + 2P.
1
Expert's answer
2020-11-06T10:08:25-0500

In equilibrium Qd = Qs, so:

60 − 0.5P = −2 + 2P,

2.5P = 62,

P = 24.8,

Q = -2 + 2×24.8 = 47.6 units.